Currently, it is known that 85 percent of the population own an automobile and 15 percent do not. Based upon past experience, in a year 1¼ of the nonowners will become owners, and 1/1000 of one percent of the owners will become nonowners.


1. What are the expcted percentages of owners and nonowners for next year?


2. What is the expected steady state?

 


1
Expert's answer
2021-03-18T08:04:44-0400

Solution:

Currently, the population having an automobile "=85\%"

And the population not having an automobile "=15\%"

In a year, non-owners who become owners "=1\frac 14\%=1.25\%"

And owners who become non-owners"=\frac 1{1000}\%=0.001\%"

(1):   For next year:

Expected "\%" of owners "= 85\% + 15\% \times 1.25\% = 85.1875\%"

Expected "\%" of non-owners "= 15\% + 85\% \times 0.001\%=15.00085\%"

(2): Expected steady state:

Of owners"=85\%(1+15\%\times1.25\%)^n"

Of non-owners"=15\%(1+85\%\times 0.001\%)^n"

where "n" is the number of years.